# Graphical tools for detection and control of selection bias with multiple exposures and samples

### v1

@@ 3.1 Within-sample identification of causal effects via separability
Figure 6 depicts a $b$-SWIG in which the isolation conditions fail due to the presence of a confounder $M^{b}$ of $S^{b}$ and $Y^{b}$ that itself is a descendant of $b$. We show in Appendix A.1 that the $b$-SWIG can be used for analyzing selection bias as a purely notational tool, without assuming the existence of hypothetically separable components of exposure. Specifically, an independence relationship holds for all finest fully randomized causally interpretable structured tree graphs (FFRCISTGs, Robins, (1986); Shpitser et al., (2022)) $\mathcal{P}$ faithful to $\mathcal{G}^{a}$ if and only if it holds for all FFRCISTGs faithful to $\mathcal{G}_{\textrm{ex}}^{b}$ and is expressible in terms of only variables in $\mathcal{P}$.
@@ A.1 Relationship between $a$-SWIGs and $b$-SWIGS
Expression of the $b$-SWIG in the abbreviated form used in this work helps to prevent errors arising in the second step by removing $B$ and $C$ from the graphs, and using $\mathcal{G}_{\textrm{ex}}^{b}$ instead of $\mathcal{G}_{\textrm{ex}}^{bc}$ prevents us from dealing with counterfactuals with $b\neq c$. Below, we provide a formal argument that the above procedure can be treated as purely notational.
@@
Now let $\mathcal{Q}$ be a FFRCISTG compatible with the expanded DAG $\mathcal{G}_{\textrm{ex}}$, with the additional determinism assumption $B^{a}\equiv C^{a}\equiv a$, which implies $A\equiv B\equiv C$ via consistency. By construction of $\mathcal{G}_{\textrm{ex}}$, each variable can have at most one of $B$ or $C$ as a parent (though variables can have both as ancestors). Supposing $V_{i}$ is a child of $B$, the recursive substitution formula yields $V_{i}^{a=\beta}=V_{i}^{b=B^{a=\beta}}=V_{i}^{b=\beta}$, and similarly $V_{i}^{a=\beta}=V_{i}^{c=\beta}$ if $V_{i}$ is a child of $C$. Thus the one-step-ahead potential outcomes in $\mathcal{P}$ for variables that are children of $A$ on $\mathcal{G}$ retain their interpretation in $\mathcal{Q}$, as do all potential outcomes under the intervention setting $A=a$ by the recursive substitution formula. In particular, letting $\boldsymbol{V}_{1}$ be variables in $\mathcal{P}$ that are descendants of $A$ and $\boldsymbol{V}_{0}$ be all other variables in $\mathcal{P}$ except $A$, the independencies among $(\boldsymbol{V}_{0},A,\boldsymbol{V}_{1}^{a})$ are the same in $\mathcal{P}$ and $\mathcal{Q}$. Thus the statements not involving $B$, $C$, or interventions on either about a FFRCISTG provable under the DAG $\mathcal{G}_{\textrm{ex}}$ plus $B^{a}\equiv C^{a}=a$ are exactly those provable under the DAG $\mathcal{G}$, though potentially with differing difficulties of proofs.

### v2

@@ 3.1 Within-sample identification of causal effects via separability
Figure 6 depicts a $b$-SWIG in which the isolation conditions fail due to the presence of a confounder $M^{b}$ of $S^{b}$ and $Y^{b}$ that itself is a descendant of $b$. We discuss with examples in Appendix A.1 the additional assumptions implied by a $b$-SWIG compared to the original SWIG. Specifically, if some set of variables $\boldsymbol{W}$ consists of descendants of $B$ but not $C$, and some set of variables $\boldsymbol{X}$ consists of descendants of $C$ but not $B$, then the $b$-SWIG makes the assumptions that $\boldsymbol{W}^{a=1}\perp\!\!\!\perp\boldsymbol{X}^{a=0}$ and $\boldsymbol{W}^{a=0}\perp\!\!\!\perp\boldsymbol{X}^{a=1}$. These assumptions (and possibly many more) are already made by non-parametric structural equation models with independent errors (NPSEM-IE) associated with the underlying DAG $\mathcal{G}$, but are not made by the finest fully randomized causally interpretable structured tree graph (FFRCISTG) associated with the $a$-SWIG $\mathcal{G}^{a}$ (Robins,, 1986; Shpitser et al.,, 2022). They are, however, made by the FFRCISTG associated with the $(b,c)$-SWIG $\mathcal{G}_{\text{ex}}^{bc}$ under the determinism assumption $a\equiv B^{a}\equiv C^{a}$, and are not cross-world if $B$ and $C$ correspond to real components of $A$ that can be intervened on separately.
@@ A.1 Relationship between $a$-SWIGs and $b$-SWIGS
Expression of the $b$-SWIG in the abbreviated form used in this work helps to prevent errors arising in the second step by removing $B$ and $C$ from the graphs, and using $\mathcal{G}_{\textrm{ex}}^{b}$ instead of $\mathcal{G}_{\textrm{ex}}^{bc}$ prevents us from dealing with counterfactuals with $b\neq c$. Below, we provide formal details of the relationship between $a$-SWIGs and $b$-SWIGs under two common causal models.
@@
Now let $\mathcal{Q}$ be a FFRCISTG compatible with the expanded DAG $\mathcal{G}_{\textrm{ex}}$, with the additional determinism assumption $B^{a}\equiv C^{a}\equiv a$, which implies $A\equiv B\equiv C$ via consistency. By construction of $\mathcal{G}_{\textrm{ex}}$, each variable can have at most one of $B$ or $C$ as a parent (though variables can have both as ancestors). Supposing $V_{i}$ is a child of $B$, the recursive substitution formula yields $V_{i}^{a=\beta}=V_{i}^{b=B^{a=\beta}}=V_{i}^{b=\beta}$, and similarly $V_{i}^{a=\beta}=V_{i}^{c=\beta}$ if $V_{i}$ is a child of $C$. Thus the one-step-ahead potential outcomes in $\mathcal{P}$ for variables that are children of $A$ on $\mathcal{G}$ retain their interpretation in $\mathcal{Q}$, as do all potential outcomes under the intervention setting $A=a$ by the recursive substitution formula. However, positing that $B$ and $C$ exist as components of $A$ on which we could intervene separately means that the FFRCISTG associated with the expanded graph assumes independence between sets of potential outcomes that are descendants of one but not the other. That is, if $\boldsymbol{W}$ consists of descendants of $B$ but not of $C$, and $\boldsymbol{X}$ consists of descendants of $C$ but not of $B$, then we assume $\boldsymbol{W}^{b}\perp\!\!\!\perp\boldsymbol{X}^{c}$, and due to determinism and recursive substitution, $\boldsymbol{W}^{a=1}\perp\!\!\!\perp\boldsymbol{X}^{a=0}$ and $\boldsymbol{W}^{a=0}\perp\!\!\!\perp\boldsymbol{X}^{a=1}$. These assumptions (and possibly many more) are made by the non-parametric structural equation model with independent errors (NPSEM-IE) for both the original and expanded graphs. Two concrete examples are given below, along with a discussion of some implications.
[…]
Figure A.1: (Left) A single world intervention graph (SWIG) $\mathcal{G}^{a}$ representing intervention on $A$. (Right) A SWIG $\mathcal{G}_{\text{ex}}^{abc}$ representing the possibility of intervening also on components $B$ and $C$ of $A$ separately.
Figure A.1 depicts an example of an $a$-SWIG $\mathcal{G}^{a}$ for intervening on exposure $A$, and an expanded SWIG $\mathcal{G}_{\text{ex}}^{abc}$ corresponding to intervening also on the components $B$ and $C$ of $A$ separately. The FFRCISTG associated with $\mathcal{G}^{a}$ makes the following independence assumptions:
$A$ $\perp\!\!\!\perp Y^{a=0}\perp\!\!\!\perp S^{a=0},$ (A.3)
$A$ $\perp\!\!\!\perp Y^{a=1}\perp\!\!\!\perp S^{a=1},$
so that the variables on each line are mutually independent. The FFRCISTG associated with $\mathcal{G}_{\text{ex}}^{abc}$ makes the larger set of assumptions:
$A\perp\!\!\!\perp B^{a}\perp\!\!\!\perp Y^{b}\perp\!\!\!\perp C^{a}\perp\!\!\!\perp S^{c}\text{ for }a,b,c\in\{0,1\}.$ (A.4)
Under the determinism restriction $a\equiv B^{a}\equiv C^{a}$, the above assumptions become
$A\perp\!\!\!\perp Y^{b=0}\perp\!\!\!\perp S^{c=0},$ (A.5)
$A\perp\!\!\!\perp Y^{b=0}\perp\!\!\!\perp S^{c=1},$
$A\perp\!\!\!\perp Y^{b=1}\perp\!\!\!\perp S^{c=0},$
$A\perp\!\!\!\perp Y^{b=1}\perp\!\!\!\perp S^{c=1},$
and further by recursive substitution the independence assumptions among variables in $\mathcal{G}^{a}$ (i.e., in terms of intervention on $A$) become
$A\perp\!\!\!\perp Y^{a=0}\perp\!\!\!\perp S^{a=0},$ (A.6)
$A\perp\!\!\!\perp Y^{a=0}\perp\!\!\!\perp S^{a=1},$
$A\perp\!\!\!\perp Y^{a=1}\perp\!\!\!\perp S^{a=0},$
$A\perp\!\!\!\perp Y^{a=1}\perp\!\!\!\perp S^{a=1}.$
Note that the first and last assumptions are the same as those made by the FFRCISTG associated with $\mathcal{G}^{a}$. The second and third assumptions include additional independencies $Y^{a=0}\perp\!\!\!\perp S^{a=1}$ and $Y^{a=1}\perp\!\!\!\perp S^{a=0}$ that are cross-world if we cannot define and intervene on the components $B$ and $C$ separately, but are not cross-world if we can do so.
[…]
Figure A.2: A fully-intervened expanded SWIG $\mathcal{G}_{\text{ex}}^{abm}$ for which the associated FFRCISTGs agrees with the FFRCISTG associated with the unexpanded SWIG $\mathcal{G}^{am}$ when $a\equiv B^{a}\equiv C^{a}$ but not the NPSEM-IE associated with the underlying unexpanded DAG $\mathcal{G}$.
In the above example, the independencies among $A$ and potential outcomes induced by intervention on it are a strict superset of those assumed by the FFRCISTG associated with $\mathcal{G}^{a}$, and exactly those assumed by the NPSEM-IE associated with the underlying DAG $\mathcal{G}$. However, neither is the case in general. Figure A.2 depicts a case where neither relationship holds. Because $C$ has no descendants in $\mathcal{G}_{\text{ex}}$, the FFRCISTG associated with $\mathcal{G}_{\text{ex}}^{abm}$ does not make any additional assumptions about variables induced by intervention on variables in $\mathcal{G}^{am}$ that the FFRCISTG associated with $\mathcal{G}^{am}$ does not make. However, the NPSEM-IE associated with $\mathcal{G}$ makes the assumptions $Y^{a=0}\perp\!\!\!\perp S^{a=1}$ and $Y^{a=1}\perp\!\!\!\perp Y^{a=0}$. In contrast to the possibly cross-world assumptions made by the FFRCISTG associated with the expanded SWIG in the previous example, these assumptions are necessarily cross-world in the sense that, even if one could intervene on any components of $A$, and even on any components of $M$ separately, one still could not observe $Y^{a=1}$ and $S^{a=0}$ together (nor $Y^{a=0}$ and $S^{a=1}$ together). The salient property of the system depicted is that $M$ is a common cause of $Y$ and $S$ that is also causally affected by $A$. It is sometimes said that to observe both $Y^{a=0}$ and $S^{a=1}$, $M$ would need to be a recanting witness (Avin et al.,, 2005) in that it would need to take the value $M^{a=0}$ to observe $Y^{a=0}$ and $M^{a=1}$ to observe $S^{a=1}$.